Ivan Industries (II) has a debt-to-equity ratio of 1.4, a corporate tax rate of 30%, pays...

  1. Ivan Industries (II) has a debt-to-equity ratio of 1.4, a corporate tax rate of 30%, pays 4% interest on its debt and has a required rate of return on equity of 12%. What is II’s WACC? How much does the debt tax shield reduce II’s WACC? What is the required rate of return on firm assets?
Answer #1

WACC = (1.4/2.4)(0.04)(1 - 0.30) + (1/2.4)(0.12)

WACC = 6.63%

Reduction in WACC by interest tax shield = (1.4/2.4)(0.04)(0.30) = 0.70%